OpenStudy (calculusxy):

Initially, Greg had a total of 60 records and cassettes in his music collection. He then sold 1/8 of his cassettes and 1/2 of his records. If the number of records he sold is twice the number of cassettes he sold, how many records did he sell? @satellite73

OpenStudy (anonymous):

i hate word problems

OpenStudy (calculusxy):

Me too.

OpenStudy (anonymous):

we need a variable, lets say \(x\) is the number of records, so \(60-x\) is the number of cassettes

OpenStudy (anonymous):

the number of records sold is \(\frac{1}{2}x\) and the number of cassettes sold is \(\frac{1}{8}(60-x)\)

OpenStudy (anonymous):

since the number of records sold is twice the number of cassettes sold, we know \[\frac{2}{8}(60-x)=\frac{1}{2}x\] and we can solve for \(x\) by which i think i mean " you can solve for \(x\) i would multiply both sides by 4 first

OpenStudy (calculusxy):

Would that be 120/8 - 2/8x=1/2x?

OpenStudy (calculusxy):

@mathstudent55

OpenStudy (anonymous):

multiply by 4 first and start with \[60-x=2x\] then it will be easy

OpenStudy (anonymous):

\[\frac{2}{8}(60-x)=\frac{1}{2}x\\ \frac{1}{4}(60-x)=\frac{1}{2}x\\ 4\times \frac{1}{4}(60-x)=4\times \frac{1}{2}x\\ 60-x=2x\] etc

OpenStudy (anonymous):

when was this question written, 1976??